# FocalCodec: Low-Bitrate Speech Coding via Focal Modulation Networks

### v1

@@ Abstract
Large language models have revolutionized natural language processing through self-supervised pretraining on massive datasets. Inspired by this success, researchers have explored adapting these methods to speech by discretizing continuous audio into tokens using neural audio codecs. However, existing approaches face limitations, including high bitrates, the loss of either semantic or acoustic information, and the reliance on multi-codebook designs when trying to capture both, which increases architectural complexity for downstream tasks. To address these challenges, we introduce FocalCodec, an efficient low-bitrate codec based on focal modulation that utilizes a single binary codebook to compress speech between 0.16 and 0.65 kbps. FocalCodec delivers competitive performance in speech resynthesis and voice conversion at lower bitrates than the current state-of-the-art, while effectively handling multilingual speech and noisy environments. Evaluation on downstream tasks shows that FocalCodec successfully preserves sufficient semantic and acoustic information, while also being well-suited for generative modeling. Demo samples, code and checkpoints are available at https://lucadellalib.github.io/focalcodec-web/.
Speech coding, Discrete representations, Vector quantization, Generative models
@@ 1 Introduction
Recent advancements in large language models (OpenAI et al., 2023; Chowdhery et al., 2024; Jiang et al., 2024a; Grattafiori et al., 2024) have led to significant progress in natural language processing, enabling breakthroughs in tasks such as summarization, translation, question answering, code generation, and retrieval. Building on this success, the research community has extended these methods to other modalities, with speech emerging as a major area of interest. The impressive performance of text-conditioned audio and speech generation models (Borsos et al., 2023; Copet et al., 2023; Kreuk et al., 2023; Wang et al., 2023; Kim et al., 2024), along with recent speech language models (Zhang et al., 2023; Hassid et al., 2023; Défossez et al., 2024; Nguyen et al., 2024), highlights the potential of token-based approaches for speech processing.
@@
Despite recent progress, current codecs still face several challenges. Acoustic codecs (Défossez et al., 2023; Kumar et al., 2023; Ji et al., 2024; Xin et al., 2024) achieve high-quality reconstruction but often rely on multiple codebooks, adding complexity to the design of downstream models. Additionally, they typically lack strong semantic representations. Hybrid codecs (Zhang et al., 2024; Liu et al., 2024; Défossez et al., 2024; Parker et al., 2024) aim to combine both acoustic and semantic information while maintaining high-quality resynthesis. Still, they often depend on complex multi-codebook designs, explicit disentanglement, distillation losses, or supervised fine-tuning. Single-codebook designs (Li et al., 2024; Guo et al., 2024; Ji et al., 2024; Xin et al., 2024; Wu et al., 2024a) offer a simpler architecture but struggle to balance compression while maintaining both reconstruction quality and effective representations for downstream tasks, especially at low bitrates. To address these limitations, we introduce FocalCodec, an efficient low-bitrate codec based on focal modulation (Yang et al., 2022) that compresses speech into the space of a single binary codebook. FocalCodec achieves competitive performance in reconstruction at lower bitrates than the current state-of-the-art under a variety of conditions while also preserving sufficient semantic and acoustic information for downstream tasks.
@@
  • We introduce FocalCodec, a novel hybrid codec featuring a unique compressor-quantizer-decompressor architecture that compresses speech using a single binary codebook at ultra-low bitrates (0.16 to 0.65 kbps).
@@
Demo samples, code and checkpoints are available on our project page [^1] https://lucadellalib.github.io/focalcodec-web/.
@@ 2 Related Work
Acoustic codecs, built on the VQ-VAE (van den Oord et al., 2017) framework, aim for high-fidelity reconstruction. Notable advancements include hierarchical RVQ (Zeghidour et al., 2021), lightweight architectures (Défossez et al., 2023), improved RVQ techniques (Kumar et al., 2023), and efficiency-driven designs (Yang et al., 2023; Ren et al., 2024; Ai et al., 2024). Recent methods explore scalar quantization (Mentzer et al., 2024; Yang et al., 2024a), Mel-spectrogram discretization (Bai et al., 2024), and novel paradigms like diffusion- and flow-based decoding (Wu et al., 2024b; Yang et al., 2024b; Pia et al., 2024). To reduce bitrate without compromising performance, multi-scale RVQ (Siuzdak et al., 2024; Qiu et al., 2024) achieves improved compression by varying frame rates in deeper quantizers. However, its hierarchical design adds complexity to downstream applications, as it requires flattening the token sequences. Single-codebook designs (Li et al., 2024; Guo et al., 2024; Ji et al., 2024; Xin et al., 2024; Wu et al., 2024a) have emerged as a simpler, efficient alternative, delivering robust performance at low bitrates. Our codec aligns with this trend, leveraging a novel focal modulation architecture and a pretrained self-supervised encoder to efficiently unify semantic and acoustic representation learning.
@@
Semantic codecs leverage self-supervised features from large models trained with contrastive objectives (Baevski et al., 2020; Hsu et al., 2021; Chen et al., 2022) and k-means clustering (Lloyd, 1982) for quantization, either from a single layer (Polyak et al., 2021; Wang et al., 2024) or multiple layers (Mousavi et al., 2024b; Shi et al., 2024). Improvements upon this paradigm include replacing k-means with RVQ (Huang et al., 2024), noise-aware (Messica & Adi, 2024) and speaker-invariant tokenization (Chang et al., 2024). While these approaches effectively capture linguistic and content-related information, they often discard much of the acoustic detail, resulting in low speaker fidelity when a vocoder is trained to resynthesize speech from these representations. Our codec adopts a self-supervised architecture similar to semantic codecs but retains acoustic detail through its novel compressor-quantizer-decompressor architecture and decoupled training strategy, ensuring high-quality reconstruction while preserving the advantages of semantic representations.
@@
Hybrid codecs combine semantic and acoustic features to balance reconstruction quality and content representation. Some methods (Ju et al., 2024; Jiang et al., 2024b; Zheng et al., 2024) employ multiple codebooks to disentangle speech into distinct subspaces, such as content, prosody, and timbre, while others (Liu et al., 2024) utilize dual encoders to separately capture content and fine-grained acoustic information. Semantic distillation (Zhang et al., 2024; Défossez et al., 2024) has also been explored to enrich the first RVQ codebook with semantic information from HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2022). More recently, Parker et al. (2024) trained a large-scale transformer-based VQ-VAE, achieving exceptional reconstruction quality at ultra-low bitrates. To enhance semantic content, they employed supervised fine-tuning on force-aligned phoneme data. Our codec also belongs to this category but instead of relying on complex multi-codebook designs with explicit disentanglement, distillation losses, or supervised fine-tuning, it is purely based on self-supervised learning. It compresses both semantic and acoustic information into a single codebook, pushing the boundaries of hybrid codec design at low bitrates.
@@ 3.1 Architecture
The proposed codec is largely based on the VQ-VAE framework but incorporates novel compressor and decompressor modules between the encoder and decoder (see Figure 1). The discriminator is used only during training and is discarded afterward.
@@
To build a hybrid codec with a simple design, without relying on distillation losses or multiple encoders, the encoder must capture both acoustic and semantic information. This ensures high-quality reconstructions and expressive tokens for training downstream models. Self-supervised models like HuBERT and WavLM retain significant acoustic information in their lower layers (Chen et al., 2022), making them suitable for hybrid codecs. For instance, Baas et al. (2023) show that a high-quality vocoder can be trained using continuous representations from layer-6 of WavLM-large. Following this approach, we use the first 6 layers of WavLM-large [^2] https://github.com/microsoft/unilm/tree/master/wavlm as our encoder. However, effective quantization is critical for approximating continuous representations with sufficient granularity. Standard k-means clustering typically fails to preserve essential acoustic details (van Niekerk et al., 2022). To address this, we introduce the novel compressor-quantizer-decompressor design based on focal modulation, which allows for granular quantization that preserves both semantic and acoustic information.
@@
The compressor maps the encoder representations to a compact, low-dimensional latent space. Optionally, it can perform temporal downsampling to further reduce the frame rate. Prior work typically relies on convolutional, recurrent, or transformer-based architectures for compression. In contrast, we introduce a novel focal downscaling module, which combines a downscaling operation with a focal block. The downscaling step applies a linear projection to compress the feature dimension, while a 1D convolution can be used instead to additionally downsample along the time dimensions. To better capture periodic patterns, we follow (Kumar et al., 2023) and apply Snake activations (Ziyin et al., 2020) after the projection.
To build a focal block, we replace the self-attention mechanism in the standard transformer block with focal modulation. Focal modulation (Yang et al., 2022) is an efficient alternative to self-attention that enables fine-to-coarse modeling and introduces useful inductive biases such as translation equivariance, explicit input dependency, time and channel specificity, and decoupled feature granularity. While originally designed for image and video processing, these properties also benefit speech modeling (Della Libera et al., 2024). Unlike self-attention, which directly computes token-wise interactions, focal modulation first aggregates the global context and then modulates local interactions based on this aggregated representation. This ensures that interactions are guided by the overall context rather than being dominated by individual tokens. Formally, focal modulation computes output representation $\mathbf{y}_{i}$ for each input feature $\mathbf{x}_{i}$ in sequence $\mathbf{x}_{1:n}$ as:
@@
where $q(\cdot)$ and $h(\cdot)$ are linear projections, and $\mathbf{z}^{\ell}_{i}\in\mathbf{z}_{1:n}^{\ell}$ and $\mathbf{g}^{\ell}_{i}\in\mathbf{g}_{1:n}^{\ell}$ are the context and gating vectors at position $i$ and focal level $\ell\in\{1,\dots,L+1\}$, with $\odot$ denoting element-wise multiplication. The context sequence $\mathbf{z}_{1:n}$ is obtained via a stack of depth-wise convolutions with exponentially increasing kernel sizes to capture dependencies from short to long range, with average pooling applied to the last level feature map to incorporate global information. Then, for each focal level, a point-wise convolution is used to compute the gating sequence $\mathbf{g}_{1:n}$. This hierarchical approach, operating at multiple granularities, makes focal modulation well-suited for processing speech features, enabling efficient and scalable representation learning in linear time while preserving long-range dependencies.
@@
To address these challenges, we employ binary spherical quantization (BSQ) (Zhao et al., 2024), originally introduced for compression of images and videos. To the best of our knowledge, this is the first successful application of binary quantization in the speech domain. BSQ belongs to the category of lookup-free quantization (LFQ) methods (Yu et al., 2024), i.e. it utilizes an implicit codebook, defined as:
@@
where $\mathrm{sign}(\cdot)$ denotes the sign function, with $\mathrm{sign}(0)$ remapped to 1 to ensure the output always lies on the hypersphere. To make the quantization differentiable, we use the straight-through estimator (Bengio et al., 2013). BSQ offers several advantages over traditional quantization methods. First, the parameter-free implicit codebook is lightweight and computationally efficient. Second, empirical evidence (Zhao et al., 2024) shows that the binary quantization bottleneck encourages high codebook utilization, even for large values of $L$. Third, the quantization error is bounded, resulting in faster convergence compared to vanilla LFQ, which does not normalize the representations. Finally, tying the codebook size to the latent dimension helps prevent performance degradation in downstream generative models when using larger codebooks (Yu et al., 2024).
@@
Most codecs use symmetric architectures, where the decoder mirrors the encoder. However, some works (Bai et al., 2024; Ji et al., 2024; Liu et al., 2024) explore asymmetric designs with larger decoders to improve reconstruction quality. In this work, we adopt an asymmetric design but prioritize the encoder, allocating $\sim\text{5}$x more parameters to it than the decoder. We argue that a strong encoder is essential for extracting robust, disentangled representations for downstream tasks. Even with a high compression rate, a smaller decoder can still generate high-quality audio while offering faster inference, which is beneficial for streaming applications. For the decoder, we choose the more efficient Vocos (Siuzdak, 2024) architecture over HiFi-GAN (Kong et al., 2020). Vocos maintains consistent feature resolution and uses inverse STFT for upsampling, minimizing aliasing and improving computational efficiency. The decoder processes features through ConvNeXt (Liu et al., 2022) blocks and projects the sequence of hidden representations to Fourier coefficients for waveform reconstruction. The final audio is synthesized using inverse STFT.
@@
Following (Kong et al., 2020), we employ a multi-period discriminator and a multi-scale discriminator. This approach slightly differs from prior work (Zeghidour et al., 2021; Défossez et al., 2023; Siuzdak, 2024; Kumar et al., 2023; Ji et al., 2024), which utilize multi-resolution and/or STFT-based discriminators in place of a multi-scale discriminator. The multi-resolution and STFT-based discriminators are particularly useful for mitigating over-smoothing artifacts in high-frequency components (Kumar et al., 2023), which are more critical for music and environmental sounds. Since our focus is on speech (i.e. medium frequency range), we stick to the simpler HiFi-GAN setup.

### v2

@@ Abstract
Large language models have revolutionized natural language processing through self-supervised pretraining on massive datasets. Inspired by this success, researchers have explored adapting these methods to speech by discretizing continuous audio into tokens using neural audio codecs. However, existing approaches face limitations, including high bitrates, the loss of either semantic or acoustic information, and the reliance on multi-codebook designs when trying to capture both, which increases architectural complexity for downstream tasks. To address these challenges, we introduce FocalCodec, an efficient low-bitrate codec based on focal modulation that utilizes a single binary codebook to compress speech between 0.16 and 0.65 kbps. FocalCodec delivers competitive performance in speech resynthesis and voice conversion at lower bitrates than the current state-of-the-art, while effectively handling multilingual speech and noisy environments. Evaluation on downstream tasks shows that FocalCodec successfully preserves sufficient semantic and acoustic information, while also being well-suited for generative modeling. Demo samples and code are available at https://lucadellalib.github.io/focalcodec-web/.
@@ 1 Introduction
Recent advancements in large language models openai2023gpt4 ; chowdhery2024palm ; jiang2024mixtralexperts ; dubey2024llama3herdmodels have led to significant progress in natural language processing, enabling breakthroughs in tasks such as summarization, translation, question answering, code generation, and retrieval. Building on this success, the research community has extended these methods to other modalities, with speech emerging as a major area of interest. The impressive performance of text-conditioned audio and speech generation models borsos2023audiolm ; copet2023musicgen ; kreuk2023audiogen ; wang2023valle ; kim2024clamtts , along with recent speech language models zhang2023speechgpt ; hassid2023textually ; defossez2024moshi ; nguyen2024spiritlminterleavedspokenwritten , highlights the potential of token-based approaches for speech processing.
@@
Despite recent progress, current codecs still face several challenges. Acoustic codecs defossez2023encodec ; kumar2023dac ; ji2024wavtokenizer ; xin2024bigcodec achieve high-quality reconstruction but often rely on multiple codebooks, adding complexity to the design of downstream models. Additionally, they typically lack strong semantic representations. Hybrid codecs zhang2024speechtokenizer ; liu2024semanticodec ; defossez2024moshi ; parker2024scaling aim to combine both acoustic and semantic information while maintaining high-quality resynthesis. Still, they often depend on complex multi-codebook designs, explicit disentanglement, distillation losses, or supervised fine-tuning. Single-codebook designs li2024singlecodec ; guo2024lscodec ; ji2024wavtokenizer ; xin2024bigcodec ; wu2024ts3codec offer a simpler architecture but struggle to balance compression while maintaining both reconstruction quality and effective representations for downstream tasks, especially at low bitrates. To address these limitations, we introduce FocalCodec, an efficient low-bitrate codec based on focal modulation yang2022focalnets that compresses speech into the space of a single binary codebook. FocalCodec achieves competitive performance in reconstruction at lower bitrates than the current state-of-the-art under a variety of conditions while also preserving sufficient semantic and acoustic information for downstream tasks.
@@
  • We introduce FocalCodec, a novel hybrid codec featuring a compressor-quantizer-decompressor architecture that compresses speech using a single binary codebook at ultra-low bitrates (0.16 to 0.65 kbps).
@@
Demo samples and code are available at https://lucadellalib.github.io/focalcodec-web/.
@@ 2 Related Work
Acoustic codecs, built on the VQ-VAE vandenord2017vqvae framework, aim for high-fidelity reconstruction. Notable advancements include hierarchical RVQ zeghidour2021soundstream , lightweight architectures defossez2023encodec , improved RVQ techniques kumar2023dac , and efficiency-driven designs yang2023hifi ; yong2024ticodec ; ai2024apcodec . Recent methods explore scalar quantization mentzer2024finite ; yang2024sqcodec , Mel-spectrogram discretization bai2024dmel , and novel paradigms like diffusion- and flow-based decoding wu2024scoredec ; yang2024ladiff ; pia2024flowmac . To reduce bitrate without compromising performance, multi-scale RVQ siuzdak2024snac ; qiu2024efficient achieves improved compression by varying frame rates in deeper quantizers. However, its hierarchical design adds complexity to downstream applications, as it requires flattening the token sequences. Single-codebook designs li2024singlecodec ; guo2024lscodec ; ji2024wavtokenizer ; xin2024bigcodec ; wu2024ts3codec have emerged as a simpler, efficient alternative, delivering robust performance at low bitrates. Our codec aligns with this trend, leveraging a novel focal modulation architecture and a pretrained self-supervised encoder to efficiently unify semantic and acoustic representation learning.
@@
Semantic codecs leverage self-supervised features from large models trained with contrastive baevski2020wav2vec2 or predictive hsu2021hubert ; chen2022wavlm objectives and k-means clustering lloyd1982kmeans for quantization, either from a single layer polyak2021discrete ; wang2024selm or multiple layers mousavi2024how ; shi2024mmm . Improvements upon this paradigm include replacing k-means with RVQ huang2024repcodec ; lajszczak2024basetts ; guo2025fireredtts ; wang2025maskgct , noise-aware messica24nast and speaker-invariant tokenization chang2024dcspins . While these approaches effectively capture linguistic and content-related information, they often discard much of the acoustic detail, resulting in low speaker fidelity when a vocoder is trained to resynthesize speech directly from these representations. To improve reconstruction quality, lajszczak2024basetts ; guo2025fireredtts ; wang2025maskgct incorporate continuous embeddings to capture prosody and speaker traits. However, this defeats the purpose of using speech tokenization for unified semantic and acoustic modeling. In contrast, our codec adopts a self-supervised architecture similar to semantic codecs but preserves both semantic content and acoustic detail through its compressor-quantizer-decompressor design and decoupled training strategy, ensuring high-quality reconstruction while preserving the advantages of semantic representations.
@@
Hybrid codecs combine semantic and acoustic features to balance reconstruction quality and content representation. Some methods ju2024facodec ; jiang2024unicodec ; zheng2024freecodec employ multiple codebooks to disentangle speech into distinct subspaces, such as content, prosody, and timbre, while others liu2024semanticodec utilize dual encoders to separately capture content and fine-grained acoustic information. Semantic distillation zhang2024speechtokenizer ; defossez2024moshi has also been explored to enrich the first RVQ codebook with semantic information from HuBERT hsu2021hubert and WavLM chen2022wavlm . More recently, Parker et al. parker2024scaling trained a large-scale transformer-based VQ-VAE, achieving exceptional reconstruction quality at ultra-low bitrates. To enhance semantic content, they employed supervised fine-tuning on force-aligned phoneme data. Our codec also belongs to this category but instead of relying on complex multi-codebook designs with explicit disentanglement, distillation losses, or supervised fine-tuning, it is purely based on self-supervised learning. It compresses both semantic and acoustic information into a single codebook, pushing the boundaries of hybrid codec design at low bitrates.
@@ 3.1 Architecture
The proposed codec is largely based on the VQ-VAE framework but incorporates compressor and decompressor modules between the encoder and decoder (see Figure˜1). The discriminator is used only during training and is discarded afterward.
@@
To build a hybrid codec with a simple design, without relying on distillation losses or multiple encoders, the encoder must capture both acoustic and semantic information. This ensures high-quality reconstructions and expressive tokens for training downstream models. Self-supervised models like HuBERT and WavLM retain significant acoustic information in their lower layers chen2022wavlm , making them suitable for hybrid codecs. For instance, Baas et al. baas2023knnvc show that a high-quality vocoder can be trained using continuous representations from layer-6 of WavLM-large. Following this approach, we use the first 6 layers of WavLM-large [^1] https://github.com/microsoft/unilm/tree/master/wavlm as our encoder. However, effective quantization is critical for approximating continuous representations with sufficient granularity. Standard k-means clustering typically fails to preserve essential acoustic details vannieakerk2022soft . To address this, we introduce a compressor-quantizer-decompressor design based on focal modulation, which allows for granular quantization that preserves both semantic and acoustic information.
@@
The compressor maps the encoder representations to a compact, low-dimensional latent space. Optionally, it can perform temporal downsampling to further reduce the frame rate. Prior work typically relies on convolutional, recurrent, or transformer-based architectures for compression. In contrast, we introduce a novel focal downscaling module, which combines a downscaling operation with a focal block. The downscaling step applies a linear projection to compress the feature dimension, while a 1D convolution can be used instead to additionally downsample along the time dimensions. To better capture periodic patterns, we follow kumar2023dac and apply Snake activations ziyin2020snake after the projection.
To build a focal block, we replace the self-attention mechanism in the standard transformer block with focal modulation. Focal modulation yang2022focalnets is an efficient alternative to self-attention that enables fine-to-coarse modeling and introduces useful inductive biases such as translation equivariance, explicit input dependency, time and channel specificity, and decoupled feature granularity. While originally designed for image and video processing, these properties also benefit speech modeling dellaliber2024focal . Unlike self-attention, which directly computes token-wise interactions, focal modulation first aggregates the global context and then modulates local interactions based on this aggregated representation. Intuitively, self-attention mixes tokens by first computing pairwise similarities and then aggregating, which can make the result sensitive to a few high-scoring neighbors. Focal modulation inverts this order: it first forms a compact, multi-scale summary of the input (local + global context) and then uses this summary to modulate each token. This ensures that interactions are guided by the overall context rather than being dominated by individual tokens, while avoiding quadratic cost.
Formally, focal modulation computes output representation $\mathbf{y}_{i}$ for each input feature $\mathbf{x}_{i}$ in sequence $\mathbf{x}_{1:n}$ as:
@@
where $q(\cdot)$ and $h(\cdot)$ are linear projections, and $\mathbf{z}^{\ell}_{i}\in\mathbf{z}_{1:n}^{\ell}$ and $\mathbf{g}^{\ell}_{i}\in\mathbf{g}_{1:n}^{\ell}$ are the context and gating vectors at position $i$ and focal level $\ell\in\{1,\dots,L+1\}$, with $\odot$ denoting element-wise multiplication. The context sequence $\mathbf{z}_{1:n}$ is obtained via a stack of depth-wise convolutions with increasing kernel sizes to capture dependencies from short to long range, with average pooling applied to the last level feature map to incorporate global information. Then, for each focal level, a point-wise convolution is used to compute the gating sequence $\mathbf{g}_{1:n}$. This hierarchical approach, operating at multiple granularities, makes focal modulation well-suited for processing speech features, enabling efficient and scalable representation learning in linear time while preserving long-range dependencies.
@@
To address these challenges, we employ binary spherical quantization (BSQ) zhao2024bsq , originally introduced for compression of images and videos. To the best of our knowledge, this is the first successful application of binary quantization in the speech domain. BSQ belongs to the category of lookup-free quantization (LFQ) methods mentzer2024finite ; yu2024lfq , i.e. it utilizes an implicit codebook, defined as:
@@
where $\mathrm{sign}(\cdot)$ denotes the sign function, with $\mathrm{sign}(0)$ remapped to 1 to ensure the output always lies on the hypersphere. To make the quantization differentiable, we use the straight-through estimator bengio2013ste . BSQ offers several advantages over traditional quantization methods. First, the parameter-free implicit codebook is lightweight and computationally efficient. Second, empirical evidence zhao2024bsq shows that the binary quantization bottleneck encourages high codebook utilization, even for large values of $L$, outperforming other lookup-free methods such as finite scalar quantization (FSQ) mentzer2024finite . Third, the quantization error is bounded, resulting in faster convergence compared to vanilla LFQ, which does not normalize the representations. Finally, tying the codebook size to the latent dimension helps prevent performance degradation in downstream generative models when using larger codebooks yu2024lfq .
@@
Most codecs use symmetric architectures, where the decoder mirrors the encoder. However, some works bai2024dmel ; ji2024wavtokenizer ; liu2024semanticodec explore asymmetric designs with larger decoders to improve reconstruction quality. In this work, we adopt an asymmetric design but prioritize the encoder, allocating $\sim\text{5}$x more parameters to it than the decoder. We argue that a strong encoder is essential for extracting robust, disentangled representations for downstream tasks. Even with a high compression rate, a smaller decoder can still generate high-quality audio while offering faster inference, which is beneficial for streaming applications. For the decoder, we choose the more efficient Vocos siuzdak2023vocos architecture over HiFi-GAN kong2020hifigan . Vocos maintains consistent feature resolution and uses inverse STFT for upsampling, minimizing aliasing and improving computational efficiency. The decoder processes features through ConvNeXt liu2022convnext blocks and projects the sequence of hidden representations to Fourier coefficients for waveform reconstruction. The final audio is synthesized using inverse STFT.
@@
Following HiFi-GAN kong2020hifigan , we employ a multi-period discriminator and a multi-scale discriminator. This approach slightly differs from prior work zeghidour2021soundstream ; defossez2023encodec ; siuzdak2023vocos ; kumar2023dac ; ji2024wavtokenizer , which utilize multi-resolution and/or STFT-based discriminators in place of a multi-scale discriminator. The multi-resolution and STFT-based discriminators are particularly useful for mitigating over-smoothing artifacts in high-frequency components kumar2023dac , which are more critical for music and environmental sounds. Since our focus is on speech (i.e. medium frequency range), we stick to the simpler HiFi-GAN setup.